# Compressed data structures for Heegaard splittings

### v1

@@ 2.2. Background on computational topology
Heegaard diagrams consist of three essential components that any candidate data structure must encode: a surface, $\Sigma_{g}$, and two multicurves, $\alpha$ and $\beta$. Starting with the surface, a graph $T=(V,E)$ embedded in $\Sigma$ is a cellular embedding if the manifold defined by cutting $\Sigma$ along $T$, $\Sigma\backslash T$, is homeomorphic to a collection of disks. We call each disk component of $\Sigma\backslash T$ a face of $T$. A cellular embedding $T$ in the surface $\Sigma$ is said to be oriented if we assign to each face an ordered list of its vertices, consistent with the orientation of $\Sigma$ as in the right-hand rule. We assume every cellular graph to be oriented. A cellular embedding $T$ is a (generalized) triangulation of $\Sigma$ if exactly three edges of $T$ bound (the closure of) each face. We define the size of a triangulation, $|T|$, as its number of faces; note that $|T|=O(|E|)$ where $|E|$ is the number of edges of the triangulation. Surface triangulations are not to be confused with the similar, but different concept of 3-manifold triangulations.
@@
Every oriented surface can be described as a data structure by an oriented triangulation. For the triangulations themselves, we will assume that they are given by the set of vertices, edges, and triangles, with bidirectional pointers between vertices and edges and between edges and the triangles. In particular, determining whether two triangles are adjacent, getting the incident vertices and edges of a triangle, and finding triangles bounded by a specific edge take constant time in this model. Other more compact data structures for triangulation, such as CGAL’s [13], can be converted into our structure in time $O(|T|)$ and at similar costs to memory. More general cellular embeddings are treated similarly.
@@
A curve $c$ is edged with respect to $T$ if it is disjoint from its faces. We represent an edged curve as a data structure by the list $E_{T}(c)=\{e\in E:|c\cap e|=e\}$, assumed ordered according to an implicit traversing orientation of $c$. Similarly, a multicurve $\gamma$ can be represented by a set of #$\#\gamma$ edge lists, one for each component, which we denote as $E_{T}(\gamma)$. Although we assume that no two components of $\gamma$ share an edge, they might share vertices. The complexity of $E_{T}(\gamma)$ is $\|E_{T}(\gamma)\|=\sum_{c\in\gamma}|E_{T}(c)|$.
@@
On the other hand, an essential multicurve $\gamma$ is called standard in $T$ if it intersects the graph only at the edges and transversely. If $\gamma$ is standard, we can represent it as a collection of words in a finite alphabet: we arbitrarily orient the edges $E$ and take $E^{\pm}=E\cup E^{-1}$ for the letters. Then for each component $c$ of $\gamma$, the word $I_{T}(c)$ in $(E^{\pm})^{*}$ is given by traversing $c$ along some arbitrary direction and, whenever we intersect an edge $e\in E$, we append $e^{i}$ to $I_{T}(c)$ where $i$ is the index of the intersection. An intersection sequence representation for $\gamma$, $I_{T}(\gamma)$, is the set $\{I_{T}(c)\}_{c\in\gamma}$. Note that while the $E_{T}(\gamma)$ is well-defined at the multicurve level, intersection sequences are defined only up to isotopies inside the cellular graph’s faces. Moreover, cyclic permutations and taking inverses of the words $I_{T}(c)$ in $I_{T}(\gamma)$ define the same isotopy class.
@@
Instead of working with explicit intersection sequences, we will often assume straight line programs (SLPs) to express $I_{T}(c)$. Formally, an SLP on the alphabet $E^{\pm}$ is a finite sequence of equations $\langle x_{1}=\text{EXPR}_{1},\dots,x_{\ell}=\text{EXPR}_{\ell}\rangle$, where the left-hand side, $x_{i}$, are symbols not in $E^{\pm}$ called variables and the expressions on the right-hand side, $\text{EXPR}_{i}\in(E^{\pm}\cup\{x_{1},\dots,x_{i-1}\})^{*}$, are called assignments. Each assignment $\text{EXPR}_{i}$ is assumed to be either a symbol from $E^{\pm}$, in which case the assignment is called simple, or of form $\text{EXPR}_{i}=x_{j_{1}}^{\pm 1}\cdot x_{j_{2}}^{\pm 1}\cdot\dots\cdot x_{j_{k}}^{\pm 1}$, where $1\leq j_{1},\dots,j_{k}<i$, in which case we say that the assignment is proper. In particular, the SLP $\langle x_{1},\dots,x_{\ell}\rangle$ represents a word $w$ in $(E^{\pm})^{*}$ if, by recursively substituting the appropriate $x_{j}$ in each proper assignment, gives $x_{\ell}$ equal to $w$. We call $\ell$ the length of the SLP and define its complexity $m=\sum_{i=1}^{\ell}|\text{EXPR}_{i}|$, where $|\text{EXPR}_{i}|$ is the number of variables and symbols in the assignment $|\text{EXPR}_{i}|$. Given any word $e_{1}\dots e_{n}$ in $(E^{\pm})^{*}$, one can trivially build an SLP of complexity $2n$, but there are instances in which SLPs can be used to give exponential gains over uncompressed representation of words [49]. Nonetheless, even in these extremely compressed cases, several useful operations can still be efficiently performed in the SLP’s complexity, as stated in the following lemma. Here and throughout, we assume that $\lg k$ is the number of bits necessary to represent the potentially signed integer $k$; in particular, $\lg k=\log k+2$ if $k$ is signed and $\log k+1$ if unsigned.
@@
$\langle x_{1},\dots,x_{\ell},x_{\ell+1}=x_{\ell}\cdot x_{\ell},\dots,x_{\ell+\lg k}=x_{\ell+\lg k-1}\cdot x_{\ell+\lg k-1},b_{0}x_{\ell}\cdot b_{1}x_{\ell+1}\cdot\dots\cdot b_{\lg k}x_{\ell+\lg k}\rangle$
@@
Now, suppose that $T=(V,E)$ is a triangulation. A standard (multi)curve $\gamma$ is normal if its intersection with each face of $T$ are arcs, called elementary arcs, that connect distinct sides of the triangle in which they are contained. Equivalently, $\gamma$ is normal if each (uncompressed) intersection sequence in $I_{T}(\gamma)$ is cyclic reduced, that is, if there are no substrings of type $ee^{-1}$ or $e^{-1}e$ in any of the cyclic permutation of the components’ sequences. When the curve $c$ is normal for a triangulation, its isotopy class is fixed by the number of times it intersects each labeled edge [48], implying that we can describe it through a vector in $N_{T}(c)$ in $\mathbb{N}^{|E|}$. We call $N_{T}(c)$ the curves’ normal coordinates, whose complexity, $\|N_{T}(c)\|$, is the number of bits necessary to store this vector, i.e., $\sum_{e\in E}\lg e(c)=\sum_{e\in E}(\log e(c)+1)$. Naturally, we define $N_{T}(\gamma)$ as the set $\{N_{T}(c)\}_{c\in\gamma}$.
@@
We could equally represent a normal multicurve $\gamma$ with a single vector $\sum_{c\in\gamma}N_{T}(\gamma)$. Similarly, it is possible and common to extend the notion of SLPs to encode more than one uncompressed word, which in our case, implies giving all components of $\gamma$ as a single SLP. Although these approaches are more parsimonious, they are less useful for our needs. Moreover, Lemma 5.1 of [18] gives an algorithm to transform, in time $O(|T|\|N_{T}(\gamma)\|)$, the normal coordinates of a multicurve given by a sum of its components’ coordinates to our expected form.
Any normal (multi)curve $\gamma$ partitions the edges of $T$ into segments called ports. The overlay graph of $\gamma$ is defined by having as vertices the union of $V$ and the intersections $E\cap\gamma$, and as edges the union of the ports and the elementary arcs of $\gamma$ (see Figure 2). The overlay graph splits each triangle $t$ of the original triangulation into two types of regions: junctions, which are adjacent to tree edges of $T$, and blocks, which are adjacent to only two. A port of the overlay graph is called redundant if it separates blocks from different triangles of the original triangulation. Following Erickson and Nayyeri, we define the street complex of $\gamma$, $S(T,\gamma)$, by deleting redundant ports and their endpoints from the overlay graph. In practice, the street complex merges blocks of the overlay graph into longer faces called streets; naturally, $S(T,\gamma)$ is a cellular graph in the surface triangulated by $T$, see Figure 2. We define the crossing sequence of a street in $S(T,\gamma)$ as the intersection sequence of a curve that crosses the street from one end to the other with respect to the original triangulation $T$. For simplicity, we assume that any port that separates two junctions is a degenerate street with a crossing sequence of length 1. The street complexes of all multicurves in this paper will have complexity $O(|T|)$ [18].
@@ 3. Data structure for Heegaard diagrams
Normal coordinates suggest a very natural encoding of Heegaard diagrams and, consequently, of closed 3-manifolds. Explicitly, we propose representing a diagram $(\Sigma_{g},\alpha,\beta)$ by a tuple $(T,E_{T}(\alpha),N_{T}(\beta))$, where $T=(V,E)$ is a triangulation, $E_{T}(\alpha)$ is an edge representation of $\alpha$, and $N_{T}(\beta)$ is a normal coordinate representation of $\beta$. We will often refer to the pair $(T,E_{T}(\alpha))$ as a marked triangulation of $\Sigma_{g}$, in which context $N_{T}(\beta)$ will be called the diagram itself. The complexity of the diagram is
@@
We note that $\|E_{T}(\alpha)\|=O(|T|)=O(\|N_{T}(\beta)\|/|\beta\cap E|)$, that is, for complexity purposes, only the last term of (3.1) matters. Nonetheless, for the sake of clarity, we will still refer to the terms of equation (3.1) arising from the marked triangulation in our analyses. All complexities will be reported under the assumption that $|\beta\cap T|=\Omega(|T|^{2})$ as, otherwise, there exist some faster trivial algorithms.
Remark 3.2.
For implementation purposes, especially in the context of Section 5, it is convenient to use the polygon representations of closed surfaces [22] to recursively define a sequence of marked triangulations $T_{g}$ of $\Sigma_{g}$, with $|T_{g}|=\Theta(g)$ and all Lickorish generators edged in $T_{g}$.
Our choices of data structures for $\alpha$ and $\beta$ allow us to represent curves with exponentially many intersection points with respect to the complexity of the diagram. Referring to the discussion at the end of Section 2.1, more intersections may lead to more complicated fundamental groups and, consequently [39][Proposition 2.6.6], more complicated 3-manifolds. In this sense, our structure is significantly more compressed than representations of Heegaard diagrams with both $\alpha$ and $\beta$ edged, which can have, at most, $E_{T}(\alpha)\times E_{T}(\beta)=O(|T|^{2})$ many intersection points. This means that, while there exist efficient algorithms to transform from the fully edged representation of Heegaard diagrams to ours—see Lemma 3.5—, the other direction is considerably harder.
Complexity gains are also noticeable when we compare our structure with triangulations of 3-manifolds. For example, while there exist algorithms to triangulate Heegaard diagrams [24], they take polynomial time on the number of intersection points between $\alpha$ and $\beta$. Moreover, as we empirically verify in Section 5, converting a Heegaard word to a diagram can be exponentially faster than converting the word to a triangulation for some particular inputs and, at least as fast, for the general case. On the downside, it is unclear how to obtain and manipulate geometric information (such as hyperbolicity) or to compute certain quantum invariants directly from Heegaard diagrams, making triangulations still preferable for these kinds of problem.
As stated in the introduction, our goal is to establish basic efficient algorithms even in this very compressed setting. For such a purpose, we will make extensive use of the next two results. The first of them is (almost) fully established by Erickson and Nayyeri [17].
@@
Part (a) corresponds to Theorems 4.7 of [18]. Part (b) follows from Theorem 4.8 of the same paper, where we simply add the variable corresponding to the desired street at the end of the SLP compressing the crossing sequences of all components of $\beta$. For part (c), use Lemma 3.4.2 in [55]. ∎
@@
A conceptually simpler algorithm for part (c) of Theorem 3.3 using only the methods of [18] is possible, but it outputs SLPs of complexity $O(m|T|)$.
For the next theorem, we might assume that a Heegaard word $\phi$ is given in power-notation form, that is, $\phi=\tau^{k_{n}}_{s_{n}}\circ\dots\circ\tau^{k_{1}}_{s_{1}}$, $s_{i}\neq s_{i+1}$ for $1\leq i<n$, and with each $k_{i}\in\mathbb{Z}$ in signed binary. A fully uncompressed Heegaard word $\phi^{\prime}=\tau^{\pm 1}_{s^{\prime}_{|\phi^{\prime}|}}\circ\dots\tau^{\pm 1}_{s^{\prime}_{1}}$, with some $s^{\prime}_{i}$ potentially equal to $s^{\prime}_{i+1}$, can be put in power-notation in time $O(|\phi^{\prime}|)$ with a unit-cost RAM model of bit-size $\lg|\phi^{\prime}|$. For convenience, we shall also assume that the Lickorish generators are given as edge curves with respect to $T$. Nonetheless, even if we suppose that $\mathcal{L}$ is given through normal coordinates or intersection sequences, the theorem still holds with similar complexity bounds by multiple applications of Theorem 4.9. Before we go to the main result, we state the following lemma.
@@
Proof of 3.5.
Suppose $\gamma$ is a simple curve. We can obtain, in time $O(|T|)$, the first barycentric subdivision $T^{\prime}$ of the triangulation $T$ [22], which we note to have size $|T^{\prime}|=6|T|$ (see blue edges and vertices of Figure 3(a)). We resort to $T^{\prime}$ to guarantee that no two vertices or edges of a same triangle are identified. One can easily update the representation $E_{T}(\gamma)$ to $E_{T^{\prime}}(\gamma)$ in time $O(|T|)$. Because $T^{\prime}$ is assumed oriented, for each edge $e_{i}\in E_{T^{\prime}}(\gamma)$, we can get, in time $O(1)$ on our data structure, the triangle $t_{i}$ to the left of $e_{\gamma}$ according to the traversing direction of $\gamma$ as defined by $E_{T^{\prime}}(\gamma)$. With the information of $t_{i}$, we can query on the other two edges $f_{i,1},f_{i,2}$ of $t_{i}$ distinct from $e_{i}$ and on the indices of the intersections, $k_{i,1},k_{i,2}=\pm 1$ of a directed arc in $t_{i}$ parallel to $e_{i}$. We also denote by $v_{i}$ the unique vertex of $t_{i}$ incident to $e_{i}$ and $e_{i+1}$ (uniqueness is guaranteed by the barycentric subdivision). We will iterate over the tuples $(f^{k_{i,1}}_{i,1},f^{k_{i,2}}_{i,2},v_{i},t_{i})$, $1\leq i\leq\|E_{T}(\gamma)\|$, to construct an uncompressed intersection sequence, $w_{\gamma}$, of a standard curve $\gamma^{\prime}$ (purple in Figure 3(a)) isotopic to $\gamma$.

### v2

@@ 2.2. Background on computational topology
Heegaard diagrams consist of three essential components that any candidate data structure must encode: a surface, $\Sigma_{g}$, and two multicurves, $\alpha$ and $\beta$. Starting with the surface, a graph $T=(V,E)$ embedded in $\Sigma$ is a cellular embedding if the manifold defined by cutting $\Sigma$ along $T$, $\T\Sigma\backslash$ T, is homeomorphic to a collection of disks. We call each disk component of $\T\Sigma\backslash$ T a face of $T$. A cellular embedding $T$ in the surface $\Sigma$ is said to be oriented if we assign to each face an ordered list of its vertices, consistent with the orientation of $\Sigma$ as in the right-hand rule. We assume every cellular embedding to be oriented. A cellular embedding $T$ is a (generalized) triangulation of $\Sigma$ if exactly three edges of $T$ bound (the closure of) each face. We define the size of a triangulation, $|T|$, as its number of faces; note that $|V|,|E|=O(|T|)$ where $|V|$ is the number of vertices of the triangulation and $|E|$ the number of edges. Surface triangulations should not be confused with the similar, but different concept of 3-manifold triangulations.
@@
Every oriented surface can be described as a data structure by an oriented triangulation. For the triangulations themselves, we will assume that they are given by the set of vertices, edges, and triangles, with bidirectional pointers between vertices and edges, and between edges and the triangles, which therefore requires a total space of $O(|T|)$. In particular, determining whether two triangles are adjacent, getting the incident vertices and edges of a triangle, and finding triangles bounded by a specific edge take constant time in this model. Other, more compact data structures for triangulation, such as CGAL’s [devillers2000triangulations], can be converted into our structure in time $O(|T|)$ and at similar costs to memory. More general cellular embeddings are treated similarly.
@@
A curve $c$ is edged with respect to $T$ if it is disjoint from its faces (red curve in Figure 3(a)). We represent an edged curve as a data structure by the list $E_{T}(c)=\{e\in E:|c\cap e|=e\}$, assumed to be ordered according to an implicit traversing orientation of $c$. Similarly, a multicurve $\gamma$ can be represented by a set of $\gamma$ edge lists, one for each component, which we denote as $E_{T}(\gamma)$. We assume that no two components can share an edge, nor can the same component be incident to a vertex more than once; nevertheless, two distinct components can share a vertex. The complexity of $E_{T}(\gamma)$ is $\|E_{T}(\gamma)\|=\sum_{c\in\gamma}|E_{T}(c)|$.
Remark 2.2.
The results of this paper, including Lemma 3.4, would still hold—sometimes with an extra parameter in the complexity—even if we allowed distinct components of edge-represented multicruves to share edges.
@@
On the other hand, an essential multicurve $\gamma$ is called standard in $T$ if it intersects the graph only at the edges and transversely (blue curve in Figure 3(a)). If $\gamma$ is standard, we can represent it as a collection of words in a finite alphabet: we arbitrarily orient the edges $E$ and take $E^{\pm}=E\cup E^{-1}$ for the letters. Then for each component $c$ of $\gamma$, the word $I_{T}(c)$ in $(E^{\pm})^{*}$ is given by traversing $c$ along some arbitrary direction and, whenever we intersect an edge $e\in E$, we append $e^{i}$ to $I_{T}(c)$ where $i$ is the index of the intersection. An intersection sequence representation for $\gamma$, $I_{T}(\gamma)$, is the set $\{I_{T}(c)\}_{c\in\gamma}$. While $E_{T}(\gamma)$ is well-defined at the multicurve level, intersection sequences are defined only up to isotopies inside the faces of the cellular embedding. Moreover, cyclic permutations and taking inverses of the words $I_{T}(c)$ in $I_{T}(\gamma)$ yield the same isotopy classes.
@@
Instead of working with explicit intersection sequences, we will often assume straight line programs (SLPs) to express $I_{T}(c)$. Formally, an SLP on the alphabet $E^{\pm}$ is a finite sequence of equations $\langle x_{1}=\text{EXPR}_{1},\dots,x_{\ell}=\text{EXPR}_{\ell}\rangle$, where the left-hand side, $x_{i}$, are symbols not in $E^{\pm}$ called variables, and the expressions on the right-hand side, $\text{EXPR}_{i}\in(E^{\pm}\cup\{x_{1},\dots,x_{i-1}\})^{*}$, are called assignments. Each assignment $\text{EXPR}_{i}$ is assumed to either be a symbol from $E^{\pm}$, in which case the assignment is called simple, or of the form $\text{EXPR}_{i}=x_{j_{1}}^{\pm 1}\cdot x_{j_{2}}^{\pm 1}\cdot\dots\cdot x_{j_{k}}^{\pm 1}$, where $1\leq j_{1},\dots,j_{k}<i$, in which case we say the assignment is proper. In particular, the SLP $\langle x_{1},\dots,x_{\ell}\rangle$ represents a word $w$ in $(E^{\pm})^{*}$ if, recursively substituting the appropriate $x_{j}$ in each proper assignment, gives $x_{\ell}$ equal to $w$. We call $\ell$ the length of the SLP and define its complexity $m=\sum_{i=1}^{\ell}|\text{EXPR}_{i}|$, where $|\text{EXPR}_{i}|$ is the number of variables and symbols in the assignment $|\text{EXPR}_{i}|$. Given any word $e_{1}\dots e_{n}$ in $(E^{\pm})^{*}$, one can trivially build an SLP of complexity $2n$, but there are instances in which SLPs can be used to give exponential gains over uncompressed representation of words [schaefer2008computing]. Nonetheless, even in these extremely compressed cases, several useful operations can still be efficiently performed with respect to the SLP’s complexity, as stated in the following lemma. Here and throughout, we assume that $\lg k$ is the number of bits necessary to represent the potentially signed integer $k$; in particular, $\lg k=\log k+2$ if $k$ is signed and $\log k+1$ if unsigned.
@@
$\langle x_{1},\dots,x_{\ell},x_{\ell+1}=x_{\ell}\cdot x_{\ell},\dots,x_{\ell+\lg k}=x_{\ell+\lg k-1}\cdot x_{\ell+\lg k-1},b_{0}x_{\ell}\cdot b_{1}x_{\ell+1}\cdot\dots\cdot x_{\ell+\lg k}\rangle$
@@
(a) A triangulation with an edged (red) and a standard (blue) curve.
(b) The three distinct isotopy classes of elementary arcs (blue) on a triangular face.
Figure 3.
Now, suppose that $T=(V,E)$ is a triangulation. A standard (multi)curve $\gamma$ is normal if it intersects each face of $T$ in arcs, called elementary arcs, that connect distinct sides of the triangle in which they are contained (Figure 3(b)). Equivalently, $\gamma$ is normal if each (uncompressed) intersection sequence in $I_{T}(\gamma)$ is cyclic reduced, that is, if there are no substrings of type $ee^{-1}$ or $e^{-1}e$ in any of the cyclic permutations of the components’ sequences. When the curve $c$ is normal for a triangulation, its isotopy class is fixed by the number of times it intersects each labeled edge [schaefer2002algorithms], which implies that we can describe it through a vector in $N_{T}(c)$ in $\mathbb{N}^{|E|}$. We call $N_{T}(c)$ the normal coordinates of the curves, whose complexity, $\|N_{T}(c)\|$, is the number of bits necessary to store this vector, i.e., $\sum_{e\in E}\lg e(c)=\sum_{e\in E}(\log e(c)+1)$. Naturally, we define $N_{T}(\gamma)$ as the set $\{N_{T}(c)\}_{c\in\gamma}$.
@@
We could equally represent a normal multicurve $\gamma$ by a single vector $\sum_{c\in\gamma}N_{T}(\gamma)$. Similarly, it is possible (and common) to extend the notion of SLPs to encode more than one uncompressed word, which, in our case, implies giving all components of $\gamma$ as a single SLP. Although these approaches are more parsimonious, they are less useful for our needs. Moreover, Lemma 5.1 of [Erickson2013] gives an algorithm to transform, in time $O(|T|\|N_{T}(\gamma)\|)$, the normal coordinates of a multicurve given by a sum of its components’ coordinates to our expected form.
Any normal (multi)curve $\gamma$ partitions the edges of $T$ into segments called ports. The overlay graph of $\gamma$ is defined by having as vertices the union of $V$ and the intersections $E\cap\gamma$, and as edges the union of the ports and the elementary arcs of $\gamma$ (see Figure 4). The overlay graph splits each triangle $t$ of the original triangulation into two types of regions: junctions, which are adjacent to three edges of $T$, and blocks, which are adjacent to only two. A port of the overlay graph is called redundant if it separates blocks from different triangles of the original triangulation. Following Erickson and Nayyeri, we define the street complex of $\gamma$, $S(T,\gamma)$, by deleting redundant ports and their endpoints from the overlay graph. In practice, the street complex merges blocks of the overlay graph into longer faces called streets; naturally, $S(T,\gamma)$ is a cellular embedding in the surface triangulated by $T$, see Figure 4. We define the crossing sequence of a street in $S(T,\gamma)$ as the intersection sequence of a curve that crosses the street from one end to the other with respect to the original triangulation $T$. For simplicity, we assume that any port that separates two junctions is a degenerate street with a crossing sequence of length 1. The street complexes of all multicurves in this paper will have complexity $O(|T|)$ [Erickson2013].
@@ 3. Data structure for Heegaard diagrams
Normal coordinates suggest a very natural encoding of Heegaard diagrams and, consequently, of closed 3-manifolds. Explicitly, we propose representing $(\Sigma_{g},\alpha,\beta)$ by a tuple $(T,E_{T}(\alpha),N_{T}(\beta))$, where $T=(V,E)$ is a triangulation, $E_{T}(\alpha)$ is an edge representation of $\alpha$, and $N_{T}(\beta)$ is a normal coordinate representation of $\beta$. We will often refer to the pair $(T,E_{T}(\alpha))$ as a marked triangulation of $\Sigma_{g}$, in which context $N_{T}(\beta)$ will be called the diagram itself. The complexity of the diagram is
@@
We note that $\|E_{T}(\alpha)\|=O(|T|)=O(\|N_{T}(\beta)\|/|\beta\cap E|)$, that is, for complexity purposes, only the last term of (3.1) matters. However, for the sake of clarity, we will still refer to the terms of equation (3.1) arising from the marked triangulation in our analyses. All complexities will be reported under the assumption that $|\beta\cap T|=\Omega(|T|^{2})$, as otherwise there are some faster trivial algorithms.
Our choices of data structures for $\alpha$ and $\beta$ allow us to represent curves with exponentially many intersection points with respect to the complexity of the diagram. Referring to the discussion at the end of Section 2.1, more intersections may lead to more complicated fundamental groups and, consequently [matveev2007algorithmic][Proposition 2.6.6], more complicated 3-manifolds. In this sense, our structure is significantly more compressed than representations of Heegaard diagrams with both $\alpha$ and $\beta$ edged, which can have, at most, $E_{T}(\alpha)\times E_{T}(\beta)=O(|T|^{2})$ intersections.
Complexity gains are also noticeable if we compare our structure with triangulations of 3-manifolds. The usual algorithm to triangulate Heegaard diagrams [he2023algorithm] takes polynomial-time in $|\alpha\cap\beta|$. Moreover, as we empirically verify in Section 5, converting a Heegaard word to a diagram can be exponentially faster than converting the word to a triangulation for some choices of input and at least as fast for the general case. On the downside, it is unclear how to obtain and manipulate geometric information (such as hyperbolicity) or to compute certain quantum invariants directly from Heegaard diagrams, making triangulations still preferable for these kinds of problems, but more on this is discussed in Section 6.
As stated in the introduction, our goal is to establish basic efficient algorithms even in this very compressed setting. For such a purpose, we will make extensive use of the next two results. The first of them is (almost) fully established by Erickson and Nayyeri [erickson2002optimally].
@@
Part (a) corresponds to Theorems 4.7 of [Erickson2013]. Part (b) follows from Theorem 4.8 of the same paper, where we simply add the variable corresponding to the desired street at the end of the SLP compressing the crossing sequences of all components of $\beta$. For Part (c), use Lemma 3.4.2 of [Stefankovic]. ∎
@@
A conceptually simpler algorithm for part (c) of Theorem 3.2 using only the methods of [Erickson2013] is possible, but it outputs SLPs of complexity $O(m|T|)$.
For the next theorem, we might assume that a Heegaard word $\phi$ is given in power-notation form, that is, $\phi=\tau^{k_{n}}_{s_{n}}\circ\dots\circ\tau^{k_{1}}_{s_{1}}$, $s_{i}\neq s_{i+1}$ for $1\leq i<n$, and with each $k_{i}\in\mathbb{Z}$ in signed binary. A fully uncompressed Heegaard word $\phi^{\prime}=\tau^{\pm 1}_{s^{\prime}_{|\phi^{\prime}|}}\circ\dots\tau^{\pm 1}_{s^{\prime}_{1}}$, with some $s^{\prime}_{i}$ potentially equal to $s^{\prime}_{i+1}$, can be put in power-notation in time $O(|\phi^{\prime}|)$ in a unit-cost RAM model of bit-size $\lg|\phi^{\prime}|$. For convenience, we shall also assume that the Lickorish generators are given as edge curves with respect to $T$. Nevertheless, even if we take them given by normal coordinates or intersection sequences, the theorem still holds with similar complexity bounds by multiple applications of Theorem 4.9. Before we move to the main result, we need the following (technical) lemma.
@@
Proof.
Suppose $\gamma$ is a simple curve. We can obtain, in time $O(|T|)$, the first barycentric subdivision $T^{\prime}$ of the triangulation $T$ [hatcher2002algebraic], which we note to have size $|T^{\prime}|=6|T|$ (see the blue edges and vertices of Figure 5(a)). We resort to $T^{\prime}$ to ensure that no two vertices or edges of the same triangle are identified. One can easily update the representation $E_{T}(\gamma)$ to $E_{T^{\prime}}(\gamma)$ in time $O(|T|)$. Because $T^{\prime}$ is assumed to be oriented, for each edge $e_{i}\in E_{T^{\prime}}(\gamma)$, we can get, in time $O(1)$ on our data structure, the triangle $t_{i}$ to the left of $e_{\gamma}$ according to the traversing direction of $\gamma$ as defined by $E_{T^{\prime}}(\gamma)$. With the information of $t_{i}$, we can query the other two edges of $t_{i}$ distinct from $e_{i}$, $f_{i,1},f_{i,2}$, and the indices of the intersections $k_{i,1},k_{i,2}=\pm 1$ of a directed arc in $t_{i}$ parallel to $e_{i}$. We also denote by $v_{i}$ the unique vertex of $t_{i}$ incident to $e_{i}$ and $e_{i+1}$ (uniqueness is guaranteed by the barycentric subdivision). We will iterate over the tuples $(f^{k_{i,1}}_{i,1},f^{k_{i,2}}_{i,2},v_{i},t_{i})$, $1\leq i\leq\|E_{T}(\gamma)\|$, to construct an uncompressed intersection sequence, $w_{\gamma}$, of a standard curve $\gamma^{\prime}$ (purple in Figure 5(a)) isotopic to $\gamma$.